if the utility function is given by u=x^2y^2 what is marginal utility?


1
Expert's answer
2021-09-06T07:07:04-0400

"U(xy) = x^2 y^2"

Marginal utility of x, MU is the first order partial derivative with respect to x and marginal utility of y, MU is the first order partial derivative with respect to y.

"MU_x =\frac{ d(x^2 y^2 ) }{ dx }= 2xy^2 \\

MUy = \frac{d(x^2 y^2 ) }{ dy }= 2x^2y"